OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph. A. –5 B. –1 C. 1 D. 5

OpenStudy (anonymous):

all you have to do is use one of the formulas for a linear equation. The easiest one is y=mx+b where x and y are coordinates or variables, m is the slope, and b is where the line crosses the y axis. For x+y=5 we rearrange the eqaution, subtract x from each side so we have y=-x+5, -x is the same as (-1)x , so the slope is -1.